# LLM-based Test-driven Interactive Code Generation: User Study and Empirical Evaluation

### v1

@@ V-B1 Identifying Task Candidates
We select task candidates from the MBPP dataset [Austin et al.(2021a)], a popular code generation benchmark, consisting of short Python functions designed to be solved by entry-level programmers. MBPP provides a natural language instruction, a set of tests, and a ground truth code implementation for each problem. We cluster functions based on problem domain, complexity and size. From each cluster we identified a set of candidate functions for which we generated a code and test completions for using a LLM. We finally selected 3 problems for the code completion tasks that best satisfied the selection criteria. These problems represent three distinct styles: MaxProduct is an algorithmic task involving dynamic programming, LowerUnderscore involves using regex for string manipulation, and FirstMissing involves a recursive binary search. The tasks are detailed in Table II.
@@ V-B2 Generating Code and Test Suggestions
To generate the code and test candidates, we give the natural language instructions from the MBPP dataset as a prompt to the OpenAI GPT-3.5-turbo chat completion endpoint with the default API parameters (temperature = 1.0). We then sample a set of 4 incorrect codes using the tests from the MBPP dataset, to identify buggy programs for each problem. We also run the set of generated tests against the set of codes to make sure at least 1 and at most 2 code suggestions are caught by the test, to restrict the number of codes that would be pruned away. Rather than manually inject bugs into the ground truth program, we choose to sample the set of buggy codes from the LLM to reflect the nature of bugs users may encounter in AI generated code.
@@ V-D Measured Variables
From the study recordings and user-submitted survey data, we collect a set of metrics on each task completed by the participants. In particular, we explore TiCoder impact across three important components of productivity: task correctness, time on task, and the self-reported effort needed to complete the task:
Time. We measured time taken to complete each task from the recordings of each participant interview. Time for each task includes time taken to evaluate any tests.
@@
Cognitive Load. After each task, we measure participants’ self-reported cognitive load via their responses to five NASA TLX questions [Hart and Staveland(1988)], a standardized approach to measuring self-reported cognitive load used widely across disciplines [Hart(2006)]. We measure the following metrics using the standard 20 point scale: mental demand, effort, perceived success, pace, and stress.
@@ V-E Evaluation of Measured Variables
We conduct an omnibus test using ANOVA to calculate the p-value of the treatment condition (the assistant used) against the measured metrics. To correct for multiple comparisons and conduct False Discovery Rate (FDR) correction[Benjamini and Hochberg(1995)] for significant pairs of conditions. We only report Omnibus p-values for pairs of conditions for which the results are statistically significant, and the direction of significance.
@@ VI RQ1: User Study Results
VI-A Impact on Task Correctness
{answerbox}
VI-B Impact on Task Time
{answerbox}
VI-C Impact on Task Induced Cognitive Load
{answerbox}
@@ VII-A Datasets
We use two Python programming datasets for our evaluation, including the sanitized version of the MBPP dataset  [Austin et al.(2021a)], dataset from Google, and the HumanEval dataset, introduced in the Codex paper [Chen et al.(2021)], to answer the research questions. MBPP consists of 427 and HumanEval of 164 examples in the format described in Sec IV-B, along with the hidden tests and reference implementations. We modify the original HumanEval dataset to remove any (non-hidden) input-output examples that are included in the docstring to avoid making the test generation task trivial.
@@ VII-B Evaluation metric
For evaluating the correctness of the generated code suggestions, we use the popular metric pass@k for evaluating the accuracy of code-generation by LLMs with respect to the hidden tests provided by each dataset [Chen et al.(2021)]. A code suggestion is correct if it passes all the hidden tests, and pass@k determines the mean expected value of choosing at least one correct code suggestion within all possible samples of size $k$, from an unranked set of suggestions. To evaluate TiCoder, we define the metric pass@k@m to denote the ranked pass@k for the code suggestions after $m\geq 1$ user queries. Recall that TiCoder outputs a ranked list of code suggestions, so pass@k@m measures if any of the top $k$ code suggestions is correct. Because pass@k@m is measuring the accuracy of a ranked list, it also provides a deterministic guarantee of accuracy at $k$ (given a set of code suggestions).
@@
MBPP text-davinci-003 49.16 86.88 68.04 75.26 77.33 77.88 78.08 73.67 80.00 81.31 81.97 82.41
code-davinci-002 48.25 89.75 68.42 76.21 79.37 81.18 81.97 73.89 81.25 84.63 85.95 86.79
CodeGen-6B 14.85 69.55 28.62 37.91 45.18 49.97 53.67 32.17 44.13 53.44 57.84 60.13
CodeGen2.5-7B 28.32 84.74 50.27 59.70 65.02 67.84 69.58 54.43 65.45 71.10 74.17 75.69
HumanEval text-davinci-003 44.13 87.80 60.70 67.54 71.41 72.18 72.81 62.63 72.90 75.86 77.71 77.36
code-davinci-002 30.49 91.49 51.66 62.65 70.30 73.11 74.37 53.52 68.12 76.37 78.65 79.38
CodeGen-6B 11.41 43.55 15.32 19.29 24.64 28.11 29.56 18.00 23.91 28.90 31.56 33.14
CodeGen2.5-7B 21.39 76.21 32.82 41.03 46.51 49.47 52.33 33.28 43.97 51.98 55.52 58.59
@@ VII-C Models and Baselines
  • OpenAI code-davinci-002 [Chen et al.(2021)] is a closed source LLM specifically designed for code completion tasks. It is based on the GPT-3 architecture containing 175B parameters. [^2] Access to this model was removed to the public by OpenAI in March 2023, but continues to be made free and available to researchers upon request.
  • OpenAI text-davinci-003 is also a closed source model of size 175B parameters, however, it is based on the GPT-3.5 architecture and InstructGPT [Ouyang et al.(2022)] and can be used for a variety of natural language tasks. Compared to other non-chat based completion models, text-davinci-003 demonstrates highly competitive performance on a number of tasks.
  • Salesforce CodeGen-6B [Nijkamp et al.(2022)] is an open source LLM, with 6B parameters, trained specifically to translate natural language instructions to code.
  • Salesforce CodeGen2.5-7B [Nijkamp et al.(2023)] is an improvement on CodeGen-6B and is slightly larger, with 7B parameters. Currently, this model is the state-of-the-art for code generation compared to other models of similar parameter size.
Our aim is to understand how TiCoder can help improve code generation accuracy across different LLMs, and not to identify the best performing model. Therefore, we use default configurations for each model, and only alter temperature. We experimented with different temperature configurations to optimize performance and diversity of generated code and test suggestions. Intuitively, a temperature closer to 1 allows LLMs to provide a more diverse set of solutions, whereas a temperature closer to 0 forces LLMs to only generate fewer solutions with the highest confidence. For all models we settle on a temperature of 0.8, as it maximizes the number of examples for which at least one correct code is produced within 100 suggestions. To account for the non-determinism of the LLM generations, for each dataset, we only query each model once to generate an initial set of 100 code and 50 test suggestions into a cache of responses. We use the same cache across all experiments that involve the specific LLM.
@@ VII-D Simulating User Response
Similar to oracle-guided inductive synthesis [Jha et al.(2010), Le et al.(2017), Jha and Seshia(2017)], we use the reference implementation $b_{p}$ as an oracle to answer if a test $(i,o)$ is consistent with the user intent, and provide the expected output $b_{p}(i)$ when the test output $o$ does not match the user intent (for TiCoder-Output). In other words, we assume that the intent of the user is precisely captured by the semantics of the (hidden) reference implementation. However, this models an idealized user interaction because, in practice, users may sometimes be unable to answer a test query in a reasonable amount of time (say, when asked about the value of the 100th prime number). As observed in the results of the user study, unlike an idealized user, real participants may sometimes provide noisy input. For example, we observe that participants are more prone to making mistakes while providing test outputs that dealt with edge cases. Therefore, using the oracle as a proxy indicates that our empirical evaluation represents an upper bound on the improvement that TiCoder can have on the benchmarks with real users.
@@ VII-E Results
As expected, text-davinci-003 and code-davinci-002, the two largest models, achieve the highest baseline performance on both datasets. Overall, we observe that both TiCoder in the TiCoder-PassFail and TiCoder-Output settings significantly improve pass@1 performance, across all models. As the number of test validation queries increase from $m=1$ to $m=5$ we also observe consistent improvement in pass@1 performance. Although the improvement is most pronounced at $m=1$, compared to baseline.
For example, on MBPP, TiCoder-PassFail improves pass@1 baseline performance of text-davinci-003 from 48.25% to 68.42%, an absolute improvement of 20.17% within one user query. TiCoder-Output improves performance to 73.89%, which is an absolute pass@1 improvement of 25.64% within one user query. This increases to almost 38% with 5 queries. While smaller models achieve lower pass@1 and pass@100 accuracy, TiCoder still provides modest boosts in accuracy. For the worst performing model, CodeGen-6B on HumanEval, TiCoder-PassFail provides an absolute pass@1 improvement of 3.91% within one interaction, and TiCoder-Output provides an absolute improvement of 6.59%. Similarly, the smaller CodeGen-6B with 4 interactions outperforms the best baseline pass@1 (text-davinci-003) for MBPP, while CodeGen2.5-7B with 3 interactions outperforms the best baseline pass@1 (text-davinci-003) for HumanEval.
Finally, as expected TiCoder-Output consistently provides higher accuracy compared to TiCoder-PassFail, since the former allows users to fix the incorrect test output. TiCoder-Output achieves an average absolute improvement of 38.43% in the code generation accuracy for both datasets and across all LLMs within 5 user interactions. However, it is worth noting that TiCoder-PassFail, even with its lightweight feedback (that generalizes to richer tests or specifications), always stays within 7% of the benefits of TiCoder-Output. This demonstrates the power of LLMs to generate test cases that satisfy user intent. Besides, unlike our idealized study where a user can always provide the correct output, our user study shows that users may have difficulty providing the correct output for a test input.
@@
Key Findings  TiCoder significantly improves pass@1 performance for all studied LLMs on both benchmarks, with performance improvements increasing with every test validation interaction. Additionally, the lightweight TiCoder-PassFail scenario always stays within 7% of the performance of TiCoder-Output even in this idealized user setting.
@@ VIII-A Tests as a developer-AI disambiguation mechanism
Results of our user study indicate that using tests as an interactive mechanism to formalize user intent, and then prune and rank AI generated code suggestions, can meaningfully improve programmer performance. Results show that both Assistant 2 (TiCoder-PassFail) and Assistant 3 (TiCoder-Output) are statistically distinguishable from Assistant 1 (the control condition without TiCoder), where no interactive test case verification or code pruning is used. Participants using Assistant 2 are significantly more likely to correctly evaluate AI generated code suggestions, and report reduced task induced cognitive load, without negative impact on time to complete each task. However, compared to Assistant 2, participants made more mistakes when validating tests with Assistant 3. Further research to explore a trade-off between the approaches, in practice, should be considered. However, we find that surfacing test cases in both forms might serve as a helpful mechanism for which to reason about generated code. While it is true that correctness likely depends on the evaluation strategies used by each participant, participants that were shown test cases by the AI Assistants performed significantly better.
Recent work has shown that current developer-AI interaction workflows, as simulated by Assistant 1 in our study, have raised new issues in the way that developers write code. Results suggest a need for interaction mechanisms that can support disambiguation; a critical feature of the usability of AI-Assistants [McNutt et al.(2023), Barke et al.(2023)]. We observed how TiCoder can surface a ranked list of tests that delineate the space of possible code suggestions, providing a concrete mechanism for identifying potential ambiguities in natural language used to prompt LLMs. Furthermore, recent work has shown that developers spend significant amount of time verifying code suggestions [Mozannar et al.(2022)][Bird et al.(2022)]. While we do not observe a statistically significant impact on the time taken to evaluate AI-generated code suggestions when using TiCoder, we do observe a significant reduction in the amount of cognitive effort required. We hypothesize that tests, which provide tangible artifacts for developers to reason about the code, can serve as a facilitating mechanism for constructing mental models of code functionality. Ultimately, this supports developers in the task of code evaluation. Future work should explore this in more detail.

### v2

@@ V-B1 Identifying Task Candidates
We select task candidates from the MBPP dataset [38], a popular code generation benchmark, consisting of short Python functions designed to be solved by entry-level programmers. MBPP provides a natural language instruction, a set of tests, and a ground truth code implementation for each problem. We cluster functions from MBPP based on problem domain, complexity as measured by cyclomatic complexity and size of the function in terms of lines of code. From each cluster we identified a set of candidate functions for which we generated a code and test completions for using a LLM. We finally selected 3 problems for the code completion tasks that best satisfied the selection criteria. These problems represent three distinct styles: MaxProduct is an algorithmic task involving dynamic programming, LowerUnderscore involves using regex for string manipulation, and FirstMissing involves a recursive binary search. The tasks are detailed in Table II.
@@ V-B2 Generating Code and Test Suggestions
To generate the code and test candidates, we give the natural language instructions from the MBPP dataset as a prompt to the OpenAI GPT-3.5-turbo chat completion endpoint with the default API parameters (temperature = 1.0). We then sample a set of 5 incorrect codes using the tests from the MBPP dataset, to identify buggy programs. We also run the set of generated tests against the set of codes to make sure at least 1 and at most 2 code suggestions are caught by the test, to restrict the number of codes that would be pruned away. Rather than manually inject bugs into the ground truth program, we choose to sample the set of buggy codes from the LLM to reflect the nature of bugs users may encounter in AI generated code.
@@ V-C Study Protocol
We employ a Latin Square Design to systematically vary the pairing of tasks and AI assistants. Each participant completes three tasks (T1, T2, and T3), each with a different AI assistant (Assistant 1, Assistant 2, Assistant 3). The order in which the participants use the AI assistants for each task is randomized to account for learning effects. This design ensures that each AI assistant is used an equal number of times across all tasks and positions, thus balancing potential order effects and providing a robust comparison of the AI assistants
@@ V-D Measured Variables
From the study recordings and user-submitted survey data, we collect a set of metrics on each task completed by the participants:
Time. We measured time taken to complete each task from the recordings of each participant interview. Time for each task includes time taken to evaluate any tests. We measure time on task to determine if the TiCoder workflow adds significant time overhead due to validation of tests, as compared to the control condition.
@@
Cognitive Load. The TiCoder workflow aims to improve code generation accuracy and reduce the number of candidate code snippets that a user needs to review, ultimately reducing the cognitive effort of code evaluation. However, TiCoder may also add additional cognitive effort, stemming from the effort required to validate tests. To test the impact of the workflow on cognitive effort, after each task we measure participants’ self-reported cognitive load via their responses to five NASA TLX questions [39], a standardized approach to measuring self-reported cognitive load used widely across disciplines [40]. We measure the following metrics using the standard 20 point scale: mental demand, effort, perceived success, pace, and stress.
@@ V-E Evaluation of Measured Variables
We conduct an omnibus test using ANOVA to calculate the p-value of the treatment condition (the assistant used) against the measured metrics. To correct for multiple comparisons and conduct False Discovery Rate (FDR) correction[41] for significant pairs of conditions. We only report Omnibus p-values for pairs of conditions for which the results are statistically significant, and the direction of significance. We chose mixed-effects models to account for individual variability (participants) and hierarchical data structures (task treatment pairs), using ANOVA for omnibus to test the significance of fixed effects (the assistant used), and conducted FDR to provide a comprehensive assessment of treatment effects while controlling for Type I errors.
@@ VI RQ1: User Study Results
VI-A Impact on Task Correctness
{answerbox}
VI-B Impact on Task Time
{answerbox}
VI-C Impact on Task Induced Cognitive Load
{answerbox}
@@ VII-A Datasets
We use two Python programming datasets for our evaluation, including the sanitized version of the MBPP dataset  [38], dataset from Google, and the HumanEval dataset, introduced in the Codex paper [1], to answer the research questions. MBPP consists of 427 and HumanEval of 164 examples in the format described in Sec IV-B, along with the hidden tests and reference implementations. We modify the original HumanEval dataset to remove any (non-hidden) input-output examples that are included in the docstring to avoid making the test generation task trivial.
@@ VII-B Evaluation metric
For evaluating the correctness of the generated code suggestions, we use the popular metric pass@k for evaluating the accuracy of code-generation by LLMs with respect to the hidden tests provided by each dataset [1]. A code suggestion is correct if it passes all the hidden tests, and pass@k determines the mean expected value of an arbitrary sample of size $k$ to contain at least one correct solution. To evaluate TiCoder, we define the metric pass@k@m to denote the ranked pass@k for the code suggestions after $m\geq 1$ user queries. Recall that TiCoder outputs a ranked list of code suggestions, so pass@k@m measures if any of the top $k$ code suggestions is correct. Given that the list of code suggestions from TiCoder are ordered, our metric pass@k@m is not a statistical measure (unlike pass@k which measures the statistical odds of any sample of size $k$ containing a correct code solution), but deterministically check if any one of the top $k$ ranked code suggestions is correct.
@@
MBPP text-davinci-003 49.16 86.88 68.04 75.26 77.33 77.88 78.08 77.00 82.20 83.38 83.59 83.75
code-davinci-002 48.25 89.75 68.42 76.21 79.37 81.18 81.97 76.97 85.51 87.10 87.56 87.71
CodeGen-6B 14.85 69.55 28.62 37.91 45.18 49.97 53.67 39.64 54.01 62.34 65.30 66.56
CodeGen2.5-7B 28.32 84.74 50.27 59.70 65.02 67.84 69.58 62.78 73.84 78.10 79.52 80.42
GPT-3.5-turbo 61.91 84.77 75.12 77.31 78.21 79.04 79.03 78.24 80.86 81.06 81.76 81.78
GPT-4-turbo 69.80 86.88 80.71 81.90 82.38 83.20 83.12 83.91 84.63 84.97 85.65 85.65
GPT-4-32k 67.13 87.35 81.56 82.62 82.97 83.11 83.70 84.78 85.28 85.31 85.40 85.79
HumanEval text-davinci-003 44.13 87.80 60.70 67.54 71.41 72.18 72.81 65.02 75.48 78.04 79.34 80.18
code-davinci-002 30.49 91.49 51.66 62.65 70.30 73.11 74.37 57.25 74.5 81.99 83.70 84.50
CodeGen-6B 11.41 43.55 15.32 19.29 24.64 28.11 29.56 26.34 32.81 39.08 44.23 48.12
CodeGen2.5-7B 21.39 76.21 32.82 41.03 46.51 49.47 52.33 35.86 51.60 61.54 65.02 68.25
GPT-3.5-turbo 59.45 89.02 73.16 76.48 77.01 77.75 79.22 73.44 76.60 78.45 78.45 79.51
GPT-4-turbo 62.62 89.63 78.36 80.42 80.92 80.84 81.34 82.22 83.17 83.42 83.62 83.84
GPT-4-32k 60.72 89.02 76.10 78.43 79.07 79.48 79.49 81.37 82.46 82.49 82.49 82.54
@@ VII-C Models and Baselines
  • OpenAI code-davinci-002, text-davinci-003 [1] is a closed source LLM specifically designed for code completion tasks. It is based on the GPT-3 architecture containing 175B parameters. [^2] Access to this model was removed to the public by OpenAI in March 2023, but continues to be made free and available to researchers upon request. text-davinci-003 is also a closed source model of size 175B parameters, however, it is based on the GPT-3.5 architecture and InstructGPT [11] and can be used for a variety of natural language tasks. Compared to other non-chat based completion models, text-davinci-003 demonstrates highly competitive performance on a number of tasks.
  • OpenAI GPT-3.5-turbo, GPT-4-turbo, GPT-4-32k The OpenAI chat models are based on the pre-trained GPT-3 model, which is fine-tuned using Reinforcement Learning with Human Feedback (RLHF) [11]. While these models are not explicitly fine-tuned for code generation, they have demonstrated strong capabilities on several related tasks [42, 43]. We use OpenAI APIs for the gpt-3.5-turbo, gpt-4-32k, and gpt-4-turbo endpoints.
  • Salesforce CodeGen-6B, CodeGen2.5-7B [3] is an open source LLM, with 6B parameters, trained specifically to translate natural language instructions to code. CodeGen2.5-7B [44] is an improvement on CodeGen-6B and is slightly larger, with 7B parameters. Currently, this model is the state-of-the-art for code generation compared to other models of similar parameter size.
Our aim is to understand how TiCoder can help improve code generation accuracy across different LLMs, and not to identify the best performing model. Therefore, we use default configurations for each model, and only alter temperature. We experimented with different temperature configurations to optimize performance and diversity of generated code and test suggestions. Intuitively, a temperature closer to 1 allows LLMs to provide a more diverse set of solutions, whereas a temperature closer to 0 forces LLMs to only generate fewer solutions with the highest confidence. Following [1, 45, 46] for all models we settle on a temperature of 0.8, as it maximizes the number of examples for which at least one correct code is produced within 100 suggestions for $k>1$ in $pass@k$. To account for the non-determinism of the LLM generations, for each dataset, we only query each model once to generate an initial set of 100 code and 50 test suggestions into a cache of responses. We use the same cache across all experiments that involve the specific LLM.
@@ VII-D Simulating User Response
Similar to oracle-guided inductive synthesis [15, 29, 28], we use the reference implementation $b_{p}$ as an oracle to answer if a test $(i,o)$ is consistent with the user intent, and provide the expected output $b_{p}(i)$ when the test output $o$ does not match the user intent (for TiCoder-Output). In other words, we assume that the intent of the user is precisely captured by the semantics of the (hidden) reference implementation. Further, if a test input crashes the reference code, we treat the user response as Undefined to model a precondition violation. However, this models an idealized user interaction because, in practice, users may sometimes be unable to answer a test query in a reasonable amount of time (say, when asked about the value of the 100th prime number). As observed in the results of the user study, unlike an idealized user, real participants may sometimes provide noisy input. For example, we observe that participants are more prone to making mistakes while providing test outputs that dealt with edge cases. Therefore, using the oracle as a proxy indicates that our empirical evaluation represents an upper bound on the improvement that TiCoder can have on the benchmarks with real users.
@@ VII-E Results
All three chat models, GPT-3.5-turbo, GPT-4-turbo and GPT-4-32k demonstrate the highest accuracy for pass@1 with comparable performance across datasets. As expected, text-davinci-003 and code-davinci-002, the two largest completion models, achieve the fourth and fifth baseline performance on both datasets. However, code-davinci-002 achieves the highest pass@100 across all models and both datasets.
Overall, we observe that both TiCoder in the TiCoder-PassFail and TiCoder-Output settings significantly improve pass@1 performance, across all models. As the number of test validation queries increase from $m=1$ to $m=5$ we also observe consistent improvement in pass@1 performance. Although the improvement is most pronounced at $m=1$, compared to baseline.
For example, on MBPP, TiCoder-PassFail improves pass@1 baseline performance of text-davinci-003 from 49.16% to 68.04%, an absolute improvement of 18.88% within one user query. TiCoder-Output improves performance to 77.00%, which is an absolute pass@1 improvement of 27.84% within one user query. This increases to 38.55% with 5 queries. While smaller models achieve lower pass@1 and pass@100 accuracy, TiCoder still provides modest boosts in accuracy. For the worst performing model, CodeGen-6B on HumanEval, TiCoder-PassFail provides an absolute pass@1 improvement of 3.91% within one interaction, and TiCoder-Output provides an absolute improvement of 14.93%.
In fact, we observe that TiCoder can boost code generation accuracy of smaller models to comparable performance of much larger SOTA LLMs. For example, after just one user interaction code-davinci-002 on MBPP achieves 68.42% accuracy in the TiCoder-PassFail setting, which is in fact higher than the pass@1 accuracy of all three SOTA chat models: GPT-4-32k, GPT-4-turbo, and GPT-3.5-turbo.
Finally, as expected TiCoder-Output consistently provides higher accuracy compared to TiCoder-PassFail, since the former allows users to fix the incorrect test output. TiCoder-Output achieves an average absolute improvement of 45.73% in the code generation accuracy for both datasets and across all LLMs within 5 user interactions. However, it is worth noting that TiCoder-PassFail, even with its lightweight feedback (that generalizes to richer tests or specifications), always stays within 9% of the benefits of TiCoder-Output. This demonstrates the power of LLMs to generate test cases that satisfy user intent.
@@
Key Findings  TiCoder significantly improves pass@1 performance for all studied LLMs on both benchmarks, with performance improvements increasing with every test validation interaction. TiCoder can boost small model performance within one user interaction, outperforming pass@1 accuracy of larger models like GPT-4-32k. Additionally, the lightweight TiCoder-PassFail scenario always stays within 9% of the performance of TiCoder-Output even in this idealized simulated user setting.
@@ VIII-A Tests as a developer-AI disambiguation mechanism
Results of our user study indicate that using tests as an interactive mechanism to formalize user intent, and then prune and rank AI generated code suggestions, can meaningfully improve programmer performance. Results show that both Assistant 2 (TiCoder-PassFail) and Assistant 3 (TiCoder-Output) are statistically distinguishable from Assistant 1 (the control condition without TiCoder), where no interactive test case verification or code pruning is used. Participants using Assistant 2 are significantly more likely to correctly evaluate AI generated code suggestions, and report reduced task induced cognitive load, without negative impact on time to complete each task. However, compared to Assistant 2, participants made more mistakes when validating tests with Assistant 3.
Further research to explore a trade-off between the approaches, in practice, should be considered. However, we find that surfacing test cases in both forms might serve as a helpful mechanism for which to reason about generated code. While it is true that correctness likely depends on the evaluation strategies used by each participant, participants that were shown test cases by the AI Assistants performed significantly better.
Recent work has shown that current developer-AI interaction workflows, as simulated by Assistant 1 in our study, have raised new issues in the way that developers write code. Results suggest a need for interaction mechanisms that can support disambiguation; a critical feature of the usability of AI-Assistants [31, 32]. We observed how TiCoder can surface a ranked list of tests that delineate the space of possible code suggestions, providing a concrete mechanism for identifying potential ambiguities in natural language used to prompt LLMs. Furthermore, recent work has shown that developers spend significant amount of time verifying code suggestions [34][14]. While we do not observe a statistically significant impact on the time taken to evaluate AI-generated code suggestions when using TiCoder, we do observe a significant reduction in the amount of cognitive effort required. We hypothesize that tests, which provide tangible artifacts for developers to reason about the code, can serve as a facilitating mechanism for constructing mental models of code functionality. Ultimately, this supports developers in the task of code evaluation. Future work should explore this in more detail.